# From Frames to Clips: Efficient Key Clip Selection for Long-Form Video Understanding

### v1

@@ Abstract
Video Large Language Models (VLMs) have achieved remarkable results on a variety of vision language tasks, yet their practical use is limited by the “needle in a haystack” problem: the massive number of visual tokens produced from raw video frames exhausts the model’s context window. Existing solutions alleviate this issue by selecting a sparse set of frames, thereby reducing token count, but such frame-wise selection discards essential temporal dynamics, leading to suboptimal reasoning about motion and event continuity. In this work we systematically explore the impact of temporal information and demonstrate that extending selection from isolated key frames to key clips - which are short, temporally coherent segments - improves video understanding. To maintain a fixed computational budget while accommodating the larger token footprint of clips, we propose an adaptive resolution strategy that dynamically balances spatial resolution and clip length, ensuring a constant token count per video. Experiments on three long-form video benchmarks demonstrate that our training-free approach, F2C, outperforms uniform sampling up to $8.1\%,5.6\%$, and $10.3\%$ on Video-MME, LongVideoBench and MLVU benchmarks, respectively. These results highlight the importance of preserving temporal coherence in frame selection and provide a practical pathway for scaling VLMs to real world video understanding applications. Project webpage is available at https://guangyusun.com/f2c.
@@ 1 Introduction
To mitigate this, most VLMs adopt uniform sampling to reduce the number of frames before encoding. While effective for short video clips, uniform sampling assumes equal importance across all frames, which is rarely the case in long and uncurated videos such as surveillance or instructional recordings. In such scenarios, uniformly sampled frames often capture irrelevant content, degrading downstream performance (Li et al., 2023; Shu et al., 2024; Liu et al., 2025).
Recent works have explored keyframe selection as a form of context management for long videos (Ataallah et al., 2024; Yu et al., 2024; Shen et al., 2024). These methods typically rank frames based on visual similarity to the query and select the most relevant ones. Although this improves efficiency, the approaches overlook temporal continuity. In long videos, sparse keyframe sampling can leave large temporal gaps. For instance, 32 frames sampled from a 30-minute video correspond to roughly 50 seconds between frames, which loses motion cues and event progression, as shown in Figure 1. This limitation raises a natural question: Can temporal information from frames adjacent to the key frames improve downstream performance in VLMs?
@@
  • We conduct a comprehensive analysis of frame selection strategies and show that selecting temporally coherent clips, rather than isolated frames, significantly improves VLM performance on long-form videos.
  • We propose a key clip selection framework, F2C, that integrates adjacent frames to restore temporal continuity, together with an adaptive resolution scheme to mitigate token overhead.
@@
  • Without any additional training, F2C outperforms uniform sampling up to 8.1%, 5.6%, and 10.3% on Video-MME (Fu et al., 2024), LongVideoBench (Wu et al., 2024b), and MLVU (Zhou et al., 2024), providing an effective and scalable solution for long-form video understanding.
Figure 1: Motivation of key clip selection. In long-form videos (e.g., 30 minutes), uniform sampling into 32 frames yields gaps of more than 50 seconds, so only a single frame may be selected from an event. As an example from Shangguan et al. (2025) shown, relying on one key frame (a) makes it impossible to answer motion-related questions such as the rotation direction of a Ping Pong ball. By extending a key frame into a short key clip (b), temporal continuity is preserved, enabling the correct reasoning.
2 Related Work
2.1 Video Large Language Model for Long-form Video
Recent advances in Video Large Language Models (VLMs) have extended the multimodal reasoning ability of LLMs to video understanding tasks such as question answering and captioning. Many works improve the backbone architecture to better handle long-form inputs: Video-LLaVA (Lin et al., 2023a), Qwen2.5-VL (Bai et al., 2025), and VideoChat2 (Li et al., 2023) adapt image-language models with temporal modules or tailored training strategies; Chat-UniVi (Jin et al., 2024) and VideoLLAMA2 (Cheng et al., 2024) refine multimodal fusion and instruction tuning; LLaVA-NeXT-QW2 (Liu et al., 2024a) and LLaVA-OneVision (Li et al., 2024) enhance vision encoders or adopt multi-granularity features; and LongVILA (Xue et al., 2024), LongVA (Zhang et al., 2024c), Video-XL (Shu et al., 2024), Video-CCAM (Fei et al., 2024), and LongVU (Shen et al., 2024) explicitly target long videos through hierarchical attention, temporal compression, or memory mechanisms. Despite these advances, most approaches still rely on uniform sampling to reduce raw videos into a fixed number of frames, overlooking redundancy and the uneven distribution of relevant information. In contrast, we focus on this preprocessing stage and propose a context management strategy that selects temporally coherent clips with adaptive resolution, offering more informative and efficient inputs without modifying the model architecture.
2.2 Training-based Frame Selection
Learning-based approaches have been proposed to select the most informative frames for VLMs, though they are typically limited to short videos due to computational cost. ATP (Buch et al., 2022) and FFS (Buch et al., 2025) train selectors end-to-end using downstream task losses, while Frame-Voyager (Yu et al., 2024) ranks frame candidates by their task loss. Hu et al. (2025) leverage a strong vision-language model to generate pseudo-labels for supervision. ViaRL (Xu et al., 2025), ReFoCUS (Lee et al., 2025) and SeViLA (Yu et al., 2023) apply reinforcement learning and self-learning to improve grounding before answering. GenS (Yao et al., 2025) and Chain-of-Frames (Ghazanfari et al., 2025) introduce datasets with keyframe annotations to enable supervised training. Although effective in controlled settings, these methods are resource-demanding to train and scale poorly to long-form videos, making training-free selection strategies more practical.
2.3 Training-free Frame Selection
Several methods have explored training-free frame selection. BOLT (Liu et al., 2025) is an early approach that leverages inverse transform sampling to improve diversity in selected frames. TCoT (Arnab et al., 2025) later utilized the long context window of Gemini-1.5-Flash (Team et al., 2024) to reason over candidate keyframes before answering. MDP3 (Sun et al., 2025) advanced this direction with a list-wise selection strategy that balances query relevance, diversity, and sequentiality. AKS (Tang et al., 2025) formulates keyframe selection as an optimization over prompt relevance and video coverage, and proposes an adaptive algorithm that selects informative keyframes to maximize retained information under fixed token budgets. Most recently, Q-Frame (Zhang et al., 2025), which is the most related to our work, incorporated dynamic resolution by ranking frames into multiple resolution levels. In contrast, we rethink frame selection from a dynamic perspective and propose an adaptive strategy over key clip length rather than treating frames as isolated units.
3 Preliminary
3.1 Problem Formulation
In this work, we consider a Video Question Answering (VQA) task for a video $V$ represented with $N$ frames $V=\{f_{1},f_{2},\dots,f_{N}\}$ and an associated text query $Q$. VQA aims to generate an answer $A$ that accurately addresses the query based on the content in the given video. Our goal is that given $(V,Q)$, we want to sample critical frames to generate the answer with the accurate information.
Due to the limitation of the computation and memory, the VLM solutions for VQA usually down-sample the video into $K$ frames based on the hardware constraints. $K$ will usually be 8 to 32 frames, i.e., $K\ll N$. By default, the frames will be sampled uniformly. Selected frames are then passed to a VLM along with the text query $Q$ to get an answer. Formally, a VLM represented by $f_{\text{VLM}}$ generates an answer for the given input as
$A=f_{\text{VLM}}(\emph{Select}(V),Q).$ (1)
So far, we only select frame with its associated spatial resolution. However, temporal information can be crucial to answer certain questions (see Figure 1). We can incorporate temporal information, key clips, into our input for VLM as $\{C_{1},C_{2},\dots,C_{K}\}$, where ${C}_{i}\in\mathbb{R}^{H\times W\times L}$. In this study, we propose an adaptive approach to incorporate key clips with adaptive spatial ($H,W$) and temporal resolution ($L$).
3.2 Importance of Key Frames and Temporal Information
To investigate the impact of key frames and the temporal information in VQA, we conduct an empirical study on the Ego4D-Haystack (Ye et al., 2025) dataset, which provides human-annotated ground-truth key frames for each video containing sufficient information to answer the question. This dataset contains long-form videos, each approximately 30 minutes in duration, which requires demanding capability on the long-context understanding. We use Qwen2.5-VL-7B as the downstream VLM to evaluate performance on the VQA task under uniform sampling, ground-truth key frames, and augmented key clips, as shown in Figure 2.
As shown in Table 1, selecting key frames based on human annotations yields a substantial improvement over uniform frame sampling, underscoring the importance of constructing task-relevant visual context. Note that even with only 3-4 key frames, key frames selection can outperform uniform sampling even with 64 frames, highlighting the strong effectiveness of selecting related frames.
To further investigate the role of temporal information, we extend both uniform sampling and keyframe selection by including neighboring frames to form short, temporally coherent segments, denoted as uniform clips and key clips. For uniform sampling, each frame is augmented into a 2-frame clip, while for keyframe selection we expand to a total of 32 frames for fair comparison. This simple augmentation consistently improves VQA accuracy, showing that preserving local temporal continuity benefits multimodal reasoning for both uniform sampling and key frames.
@@ 4.1 Anchor Key Frame Selection
For relevancy, we leverage a pre-trained contrastive vision-language encoder $E$ (e.g., CLIP (Radford et al., 2021) or SigLIP (Zhai et al., 2023; Tschannen et al., 2025)) to compute the cosine similarity between each frame $f_{i}$ and the text query $Q$ as
@@
Once relevancy scores are computed, anchor frames are selected by combining relevance with diversity. Inspired by the Watershed algorithm (Haralick et al., 1987), we first identify local maxima of the similarity curve (i.e., watersheds) as an initial candidate pool. To further ensure diversity, we apply K-means clustering on the temporal indices of these candidates into $K_{\text{anchor}}$ clusters, and select within each cluster the frame with the highest similarity score as the anchor key frame. We provide a pseudo-code of the algorithm in Appendix D.2.
Figure 3: Overview of F2C. A video and text query are first processed by a vision-language encoder to select (a) anchor key frames. Each anchor is then extended into a short key clip through (b) adaptive clip length selection, which adjusts trade-off between resolution and clip length under the same budget. The constructed key clips preserve temporal continuity and, together with the query, are fed into a VLM to generate the answer.
@@ 4.2 Adaptive Clip Length Selection
4.2.1 Exploring Trade-off between Resolution and Clip Length
@@
$B_{\text{clip}}=\frac{K\cdot L\cdot(H/s)\cdot(W/s)}{Z}=\frac{K_{\text{anchor}}\cdot L}{s^{2}\cdot K}\cdot B.$ (4)
@@
Table 2: Evaluation results of VQA accuracy across different frame budgets on Video-MME, LongVideo Bench, and MLVU. Missing results are shown as “-”. means our re-implementation results. Methods with uses various resolutions, and their budgets are equivalent to the reported # frames with the original resolution.
Method Size # Frames Video-MME (w.o./w. sub.) LongVideo Bench MLVU
@@
VLM w/ Uniform Sampling:
Video-LLaVA (Lin et al., 2023a) 7B 8 45.3 / 46.1 38.0 / 40.7 36.2 / 38.1 39.9 / 41.6 39.1 47.3
Qwen-VL (Bai et al., 2023) 7B 8 46.9 / 47.3 38.7 / 40.4 37.8 / 37.9 41.1 / 41.9 - -
VideoChat2 (Li et al., 2023) 7B 8 48.3 / 52.8 37.0 / 39.4 33.2 / 39.2 39.5 / 43.8 39.3 44.5
Chat-UniVi-V1.5 (Jin et al., 2024) 7B 8 45.7 / 51.2 40.3 / 44.6 35.8 / 41.8 40.6 / 45.9 - -
VideoLLaMA2 (Cheng et al., 2024) 7B 16 56.0 / - 45.4 / - 42.1 / - 47.9 / - - -
LLaVA-NeXT-QW2 (Liu et al., 2024a) 7B 8 58.0 / - 47.0 / - 43.4 / - 49.5 / - - -
LongVILA (Xue et al., 2024) 8B 128 60.2 / - 48.2 / - 38.8 / - 49.2 / - - -
LongVA (Zhang et al., 2024c) 7B 128 61.1 / 61.6 50.4 / 53.6 46.2 / 47.6 52.6 / 54.3 - -
Video-XL (Shu et al., 2024) 7B 128/256 64.0 / 67.4 53.2 / 60.7 49.2 / 54.9 55.5 / 61.0 - 64.9
LLaVA-OneVision (Li et al., 2024) 7B * 64.0 / 67.4 53.2 / 60.7 49.2 / 54.9 58.2 / - 56.3 64.7
Video-CCAM (Fei et al., 2024) 9B 96 61.9 / 63.1 49.2 / 52.3 39.6 / 42.4 50.3 / 52.6 - 58.5
LongVU (Shen et al., 2024) 7B 1fps 64.7 / - 58.2 / - 59.5 / - 60.9 / - - 65.4
Training-based Selection:
Frame-Voyager (Yu et al., 2024) 7B 8 67.3 / - 56.3 / - 48.9 / - 57.5 / - - 65.6
Hu et al. (2025) 8.5B 128 69.6 / - 54.1 / - 51.9 / - 58.7 / - - -
GenS (Yao et al., 2025) 7B 54 - / - - / - - / - - / - 58.7 64.8
Training-free Selection:

### v2

# From Frames to Clips: Training-free Adaptive Key Clip Selection for Long-Form Video Understanding

@@ Abstract
Video Large Language Models (VLMs) have achieved strong performance on various vision-language tasks, yet their practical use is limited by the massive number of visual tokens produced from raw video frames, which quickly exhausts the model’s context window. Existing solutions mitigate this issue by selecting a sparse set of frames, but such frame-wise selection discards essential temporal dynamics in long-form videos, leading to suboptimal reasoning about motion and event continuity. In this work, we systematically examine the role of temporal information and show that extending selection from isolated key frames to temporally coherent key clips improves video understanding. To maintain a fixed computational budget while accommodating the larger token footprint of clips, we introduce frame resolution as a controllable factor in frame selection, enabling a trade-off between spatial resolution and clip length. Building on this idea, we propose an adaptive clip length module that dynamically balances these factors to ensure a constant token count per video. Experiments on three long-form video benchmarks demonstrate that our training-free approach, F2C, outperforms uniform sampling by up to $8.1\%$, $5.6\%$, and $10.3\%$ on Video-MME, LongVideoBench, and MLVU, respectively. These results highlight the importance of preserving temporal coherence in frame selection and provide a practical pathway for scaling VLMs to real-world video understanding applications.
@@ 1 Introduction
Figure 1: Motivation of key clip selection. In long-form videos (e.g., 30 minutes), uniform sampling into 32 frames yields gaps of more than 50 seconds, so only a single frame may be selected from an event. As an example from [shangguan2025tomatoassessingvisualtemporal] shown, relying on one key frame (a) makes it impossible to answer motion-related questions such as the rotation direction of a Ping Pong ball. By extending a key frame into a short key clip (b), temporal continuity is preserved, enabling the correct reasoning.
To mitigate this, most VLMs adopt uniform sampling to reduce the number of frames before encoding. While effective for short video clips, uniform sampling assumes equal importance across all frames, which is rarely the case in long and uncurated videos such as surveillance or instructional recordings. In such scenarios, uniformly sampled frames often capture irrelevant content, degrading downstream performance [li2023videochat, shuVideoXLExtraLongVision2024, BOLT].
Recent works have explored keyframe selection as a form of context management for long videos [ataallahGoldfishVisionLanguageUnderstanding2024a, yuFrameVoyagerLearningQuery2024, shenLongVUSpatiotemporalAdaptive2024]. These methods typically rank frames based on visual similarity to the query and select the most relevant ones. Although this improves efficiency, the approaches overlook temporal continuity. In long videos, sparse keyframe sampling can leave large temporal gaps. For instance, 32 frames sampled from a 30-minute video correspond to roughly 50 seconds between frames, which loses motion cues and event progression, as shown in Figure 1. This limitation raises a natural question: Can temporal information from frames adjacent to the key frames improve downstream performance in VLMs?
@@
  • We provide a comprehensive analysis of frame selection strategies and show that selecting temporally coherent key clips, rather than isolated frames, substantially improves VLM performance on long-form videos.
  • To manage the increased tokens introduced by clips, we propose F2C, a key clip selection framework enabling an adaptive trade-off between spatial resolution and temporal length for each clip.
@@
  • Without training, F2C consistently outperforms uniform sampling, improving performance by up to 8.1%, 5.6%, and 10.3% on Video-MME [fu2024video], LongVideoBench [wu2024longvideobench], and MLVU [zhou2024mlvu], respectively, offering an efficient and scalable solution for long-form video understanding.
@@
2 Related Work
2.1 VLM for Long-form Video
Recent advances in VLMs have extended the multimodal reasoning ability of LLMs to video understanding tasks such as question answering and captioning [song2025moviechat]. Many works improve the backbone architecture to better handle long-form inputs: Video-LLaVA [lin2023video], Qwen2.5-VL [bai2025qwen25vltechnicalreport], and VideoChat2 [li2023videochat] adapt image-language models with temporal modules or tailored training strategies; Chat-UniVi [jin2024chat] and VideoLLaMA2 [chengVideoLLaMA2Advancing2024] refine multimodal fusion and instruction tuning; LLaVA-NeXT-QW2 [liu2024llavanext], LLaVA-OneVision [liLLaVAOneVisionEasyVisual2024a], and SlowFast-LLaVA [xu2024slowfast, xu2025slowfastllava15familytokenefficientvideo] and $\infty$-Video [dos2025video] enhance vision encoders or incorporate multi-scale temporal modeling; and LongVILA [xueLongVILAScalingLongContext2024], LongVA [zhang2024longva], Video-XL [shuVideoXLExtraLongVision2024], Video-CCAM [fei2024videoccamenhancingvideolanguageunderstanding], LongVU [shenLongVUSpatiotemporalAdaptive2024], ViLAMP [vilamp], and Keye-VL [keye] explicitly target long videos through hierarchical attention, temporal compression, or improved long-context modeling. Despite these advances, most approaches still rely on uniform sampling to reduce raw videos into a fixed number of frames, overlooking redundancy and the uneven distribution of relevant information. In contrast, we focus on this preprocessing stage and propose a context management strategy that selects temporally coherent clips with adaptive resolution, providing more informative and efficient inputs without modifying the model architecture.
2.2 Training-based Frame Selection
Learning-based approaches have been proposed to select the most informative frames for VLMs, though they are typically limited to short videos due to computational cost. ATP [buch2022revisitingvideovideolanguageunderstanding] and FFS [buch2025flexible] train selectors end-to-end using downstream task losses, while Frame-Voyager [yuFrameVoyagerLearningQuery2024] ranks frame candidates by their task loss. [Hu2025] leverage a strong vision-language model to generate pseudo-labels for supervision. ViaRL [xu2025viarladaptivetemporalgrounding], ReFoCUS [lee2025refocusreinforcementguidedframeoptimization] and SeViLA [yu2023self] apply reinforcement learning and self-learning to improve grounding before answering. GenS [yao2025generative] and Chain-of-Frames [ghazanfari2025chainofframesadvancingvideounderstanding] introduce datasets with keyframe annotations to enable supervised training. Although effective in controlled settings, these methods are resource-demanding to train and scale poorly to long-form videos, making training-free selection strategies more practical.
2.3 Training-free Frame Selection
Several methods have explored training-free frame selection. BOLT [BOLT] is an approach that leverages inverse transform sampling to improve diversity in selected frames. TCoT [arnab2025temporal] and CoS [hu2025coschainofshotpromptinglong] later utilized the long context window of models like Gemini-1.5-Flash [geminiteam2024gemini15unlockingmultimodal] to reason over candidate keyframes before answering. MDP3 [sun2025mdp3trainingfreeapproachlistwise] advanced this direction with a list-wise selection strategy that balances query relevance, diversity, and sequentiality. AKS [aks] formulates keyframe selection as an optimization over prompt relevance and video coverage, and proposes an adaptive algorithm that selects informative keyframes to maximize retained information under fixed token budgets. Most recently, Q-Frame [zhang2025qframequeryawareframeselection], which is the most related to our work, incorporated dynamic resolution by ranking frames into multiple resolution levels. In contrast, we rethink frame selection from a dynamic perspective and propose an adaptive strategy over key clip length rather than treating frames as isolated units.
3 Preliminary
3.1 Problem Formulation
In this work, we consider a Video Question Answering (VQA) task for a video $V$ represented with $N$ frames $V=\{f_{1},f_{2},\dots,f_{N}\}$ and an associated text query $Q$. VQA aims to generate an answer $A$ that accurately addresses the query based on the content in the given video. Our goal is that given $(V,Q)$, we want to sample critical frames to generate the answer with accurate information.
Due to the limitations of the computation and memory, the VLM solutions for VQA usually down-sample the video into $K$ frames based on the hardware constraints. $K$ will usually be 8 to 32 frames, i.e., $K\ll N$. By default, the frames will be sampled uniformly. Selected frames are then passed to a VLM along with the text query $Q$ to get an answer. Formally, a VLM represented by $f_{\text{VLM}}$ generates an answer for the given input as
$A=f_{\text{VLM}}(\emph{Select}(V),Q).$ (1)
So far, we have only selected frames with their associated spatial resolution. However, temporal information can be crucial to answer certain questions (see Figure 1). We can incorporate temporal information, key clips, into our input for VLM as $\{C_{1},C_{2},\dots,C_{K}\}$, where ${C}_{i}\in\mathbb{R}^{H\times W\times L}$. In this study, we propose an adaptive approach to incorporate key clips with adaptive spatial ($H,W$) and temporal resolution ($L$).
Figure 3: Overview of F2C. A video and text query are first processed by a vision-language encoder to select (a) anchor key frames. Each anchor is then extended into a short key clip through (b) adaptive clip length selection, which adjusts the trade-off between resolution and clip length under the same budget. The constructed key clips preserve temporal continuity and, together with the query, are fed into a VLM to generate the answer.
3.2 Key Frames and Temporal Information
To investigate the impact of key frames and the temporal information in VQA, we conduct an empirical study on the Ego4D-Haystack [tstar] dataset, which provides human-annotated ground-truth key frames for each video containing sufficient information to answer the question. This dataset contains long-form videos, each approximately 30 minutes in duration, which requires demanding capability for long-context understanding. We use Qwen2.5-VL-7B as the downstream VLM to evaluate performance on the VQA task under uniform sampling, ground-truth key frames, and augmented key clips, as shown in Figure 2.
As shown in Table 1, selecting key frames based on human annotations yields a substantial improvement over uniform frame sampling, underscoring the importance of constructing task-relevant visual context. Note that even with only 3-4 key frames, key frame selection can outperform uniform sampling even with 64 frames, highlighting the strong effectiveness of selecting related frames.
To further investigate the role of temporal information, we extend both uniform sampling and keyframe selection by including neighboring frames to form short, temporally coherent segments, denoted as uniform clips and key clips. For uniform sampling, each frame is augmented into a 2-frame clip, while for keyframe selection, we expand to a total of 32 frames for fair comparison. This simple augmentation consistently improves VQA accuracy, showing that preserving local temporal continuity benefits multimodal reasoning for both uniform sampling and key frames.
@@ 4.1 Anchor Key Frame Selection
For relevancy, we leverage a pre-trained contrastive vision-language encoder $E$ (e.g., CLIP [radford2021learning] or SigLIP [zhai2023sigmoid, tschannen2025siglip2multilingualvisionlanguage]) to compute the cosine similarity between each frame $f_{i}$ and the text query $Q$ as
@@
Once relevancy scores are computed, anchor frames are selected by combining relevance with diversity. Inspired by the Watershed algorithm [haralick1987image], we first identify local maxima of the similarity curve (i.e., watersheds) as an initial candidate pool. To further ensure diversity, we apply K-means clustering on the temporal indices of these candidates into $K_{\text{anchor}}$ clusters, and select within each cluster the frame with the highest similarity score as the anchor key frame. We provide a pseudo-code of the algorithm in the supplementary D.2.
@@ 4.2 Adaptive Clip Length Selection
4.2.1 Trade-off between Resolution and Clip Length
@@
$B_{\text{clip}}=\frac{K_{\text{anchor}}\cdot L\cdot(H/s)\cdot(W/s)}{Z}=\frac{K_{\text{anchor}}\cdot L}{s^{2}\cdot K}\cdot B.$ (4)
@@
Table 2: Evaluation results of VQA accuracy across different frame budgets on Video-MME, LongVideoBench (LVB), and MLVU. Missing results are shown as “-”. means our re-implementation results. * means # frames as a hyperparameter. Methods with uses various resolutions, and their budgets are equivalent to the reported # frames with the original resolution.
Method Size # Frames Video-MME (w.o./w. sub.) LVB MLVU
@@
Training-based VLM w/ Uniform Sampling:
Video-LLaVA [lin2023video] 7B 8 45.3 / 46.1 38.0 / 40.7 36.2 / 38.1 39.9 / 41.6 39.1 47.3
Qwen-VL [baiQwenVLVersatileVisionLanguage2023] 7B 8 46.9 / 47.3 38.7 / 40.4 37.8 / 37.9 41.1 / 41.9 - -
VideoChat2 [li2023videochat] 7B 8 48.3 / 52.8 37.0 / 39.4 33.2 / 39.2 39.5 / 43.8 39.3 44.5
Chat-UniVi-V1.5 [jin2024chat] 7B 8 45.7 / 51.2 40.3 / 44.6 35.8 / 41.8 40.6 / 45.9 - -
VideoLLaMA2 [chengVideoLLaMA2Advancing2024] 7B 16 56.0 / - 45.4 / - 42.1 / - 47.9 / - - -
LLaVA-NeXT-QW2 [liu2024llavanext] 7B 8 58.0 / - 47.0 / - 43.4 / - 49.5 / - - -
LongVILA [xueLongVILAScalingLongContext2024] 8B 128 60.2 / - 48.2 / - 38.8 / - 49.2 / - - -
LongVA [zhang2024longva] 7B 128 61.1 / 61.6 50.4 / 53.6 46.2 / 47.6 52.6 / 54.3 - -
Video-XL [shuVideoXLExtraLongVision2024] 7B 128/256 64.0 / 67.4 53.2 / 60.7 49.2 / 54.9 55.5 / 61.0 - 64.9
LLaVA-OneVision [liLLaVAOneVisionEasyVisual2024a] 7B * 64.0 / 67.4 53.2 / 60.7 49.2 / 54.9 58.2 / - 56.3 64.7
Video-CCAM [fei2024videoccamenhancingvideolanguageunderstanding] 9B 96 61.9 / 63.1 49.2 / 52.3 39.6 / 42.4 50.3 / 52.6 - 58.5
LongVU [shenLongVUSpatiotemporalAdaptive2024] 7B 1fps 64.7 / - 58.2 / - 59.5 / - 60.9 / - - 65.4
SF-LLaVA-1.5 [xu2025slowfastllava15familytokenefficientvideo] 7B 128 - / - - / - - / - 63.9 / - 62.5 71.5
ViLAMP [vilamp] 7B 1fps - / - -/ - 57.8 / 68.1 67.5 / 73.5 61.2 -
Keye-VL-1.5 [keye] 8B 64 81.2 / 84.0 70.7 / 71.9 67.1 / 66.2 73.0 / 74.0 66.0 -
Training-based Frame Selection:
Frame-Voyager [yuFrameVoyagerLearningQuery2024] 7B 8 67.3 / - 56.3 / - 48.9 / - 57.5 / - - 65.6
Hu et al. [Hu2025] 8.5B 128 69.6 / - 54.1 / - 51.9 / - 58.7 / - - -
GenS [yao2025generative] 7B 54 - / - - / - - / - - / - 58.7 64.8
Training-free Frame Selection:
@@
Table 3: Ablation study on Video-MME.
Method Short Medium Long Overall
Uniform 67.3 55.0 48.9 57.1
Watershed 70.0 58.3 50.7 59.7
+ Key Clip (Fixed) 73.0 60.0 53.7 62.2
+ Key Clip (Adaptive) 73.8 62.2 54.1 63.4
Table 4: Impact of temporal continuity. We evaluate Baselines use 64 isolated low-resolution frames, while F2C uses 16 clips of up to 4 frames. The subscript means the performance delta compared to using 16 high-resolution frames.
Method # Frames Short Medium Long Overall
Uniform 64 71.3+4.0 60.1+5.1 51.4+2.5 61.0+3.9
Top-$k$ 64 73.6+3.4 59.9+2.5 52.3+1.5 61.9+2.4
BOLT 64 73.4+1.7 61.0+3.3 50.6+0.9 61.7+2.0
Watershed 64 72.3+2.3 59.8+1.5 53.4+2.7 61.9+2.2
[…]